Shelly is 3 years older than Michele. Four years ago the sum of their age was 67. Find the age of each person now.
Answer by macston(5194)

.
If the sum of ages 4 years ago was 67, current sum is 75
(67 + 4 years for each person)
.
S=Shelly's age; M=Michelle's age=S-3
.
S+M=75
S+S-3=75
2S=78
S=39 ANSWER 1: Shelly is 39 years old.
M=S-3=39-3=36 ANSWER 2: Michelle is 36 years old.
.
CHECK:
4 years ago, sum was 67:
(S-4)+(M-4)=67
(39-4)+(36-4)=67
35+32=67
67=67